If you have 10 cars. 2 is yellow and 4 is blue and 4 is red what percention is red?
Answer by jim_thompson5910(28715)

(4/10)*100 = 0.4*100 = 40%

So 40% of the cars are red.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim